OpenStudy (anonymous):

Show all work to solve 1/x + 3/8= 1/2. Use complete sentences to explain how to solve this equation.

OpenStudy (anonymous):

subtract both side 3/8 1/x+3/8-3/8=1/2-3/8 1/x=1/2-3/8 1/2=4/8 1/x=4/8-3/8 1/x=(4-3)/8 1/x=1/8 multiply by x both side 1=x/8 now multiply by 8 both side 8=x

OpenStudy (anonymous):

thank u